ID No Question company is randomly Let individual was cited during the last 3 years. The pmf of Y is 15 points) An individual who has automobile insurance individud s andomly seleal who has automoble insurance from a certarn selected. Let Y be the number of moving violations for which the 050.10 p(y) 0.25 0.05 (1). Find a, such that py) is a p.m.f. (2).Write out the c.d.f of Y completely (3).Suppose an individual with Y violations incurs a surcharge o expected amount of the surcharge. $( +100y). Calculate the
Answer #1

TOPIC: Probability mass function,cumulative distribution function and Expected values.

IS→ 2 3 Tata トは) | a | 0,25 | 0,1010.05| 4 0 Id pCy) is a rYnt we must have, at0.25 to. Ioto, 05리. 10 | 오 | 2 | 3 | Total

0, 85 25 ta 60 rs s o. 95. So, we gee, | F(F)-ㄑ 0.60 oko ) O 75 1 ③ We-needノ Expected amount. aharae. -E(y) + 1oo E(y). we get, E(y) =d, b(8) 0.25 + 0-10 to, 15.- o-Soand, £69 - 42 2 I1O. Ex ected amount ơf the surcharge E (x+100 2) the surcharge i = E (y) + 100 E(y*). 0.50It0 1 10 5